OpenStudy (anonymous):

Im terribly lost, can someone help? A line has a slope of −5 and contains the point (−2, 4). Which equations represent the line? Choose all answers that are correct. A. −5x + y = −6 B. 5x + y = −6 C. (y – 4) = −5(x + 2) D. (y + 4) = −5(x − 2)

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

We can use the point slope form. Which is y-y1=m(x-x1) . The point is in (x,y) form so that means you can plug in the points into the formula. Also m= -5. You'll have y-4=-5(x+2)

OpenStudy (phi):

I would try x= -2 and y=4 in each of the choices and see if I could rule out any of them. can you do that ?